OpenStudy (anonymous):

You have $12000 to invest into two accounts. One account pays 4% annual simple interest and the other pays 5%. If you earned $530 annual interest, how much did you invest in each account? Any help would be greatly appreciated! I really need help with the process.

OpenStudy (gorv):

let you invested x in one account

OpenStudy (gorv):

reamain amount in second remain amount= 12000-x

OpenStudy (gorv):

now calculate interst on each

OpenStudy (gorv):

sum of interest = 530

OpenStudy (gorv):

\[530=\frac{ x*4*1 }{ 100 }+\frac{ (12000-x)*5*1 }{ 100 }\]

OpenStudy (gorv):

interest = (amount * rate*time )

OpenStudy (gorv):

now give it a try i m here to helpu

OpenStudy (anonymous):

Ill right think you ill work it out now and ge back with you. thank you so much!

OpenStudy (gorv):

wlcm

OpenStudy (anonymous):

ok so i got 7000 for one account and 5000 for the other. does that sound right?

OpenStudy (gorv):

yepppppppppp